Please help me solve this problem:
what percent of 400 is 85.4?

Answer by jim_thompson5910(35256)

Let x = unknown percent

"what percent of 400 is 85.4" translates to the equation

(x/100)*400 = 85.4

Now solve for x


x/100 = 85.4/400

x/100 = 0.2135

x = 0.2135*100

x = 21.35


So 21.35 % of 400 is 85.4